OpenStudy (anonymous):

SOmeone please help me factor this completely. (x − 5)(x + 8)^2 − (x − 5)^2(x + 8)

OpenStudy (saifoo.khan):

(x + 8)- (x − 5)

OpenStudy (anonymous):

is that function = zero?

OpenStudy (anonymous):

no just an expression. saifoo is that as low as u can factor it

OpenStudy (saifoo.khan):

Yeah, divide this by ((x − 5)(x + 8)

OpenStudy (anonymous):

I dont understand what is the answer

OpenStudy (saifoo.khan):

(x + 8)- (x − 5)

OpenStudy (anonymous):

k

OpenStudy (saifoo.khan):

want the working?

OpenStudy (anonymous):

no i think i understand. can you help me with another one.

OpenStudy (saifoo.khan):

Sure. Post it on Left/.

OpenStudy (anonymous):

(x-5) (x+8) {(x+8) - (x-5)} =(x-5) (x+8) { (x-x+8+5)} \[= (x ^{2} + 8x - 5x -40 ) {(13)}\] \[= (x ^{2} +3x -40) (13)\] \[=13 x ^{2} + 39 x -520\]